OpenStudy (anonymous):

Hello I need help solving these system of equations. Step by step not just the answer. Thank you. x + 3y = -1 2x + 2y = 6 A(-4, 1) B(2, -1) C(4, -1) D(5, -2)

OpenStudy (jhannybean):

So lets say we wanted to eliminate x if we added both of the equations together, hwo would you do that?

OpenStudy (anonymous):

I knew I needed to eliminate a variable, I'm just not sure how to set up the equations together. 2(3y -1) + 2y=-6 ??? @Jhannybean

OpenStudy (anonymous):

Thank you for your time btw

OpenStudy (jhannybean):

First try multiplying -2 to the entire first equation. -2(x + 3y = -1) 2x + 2y = 6 _____________________________ Now rewrite both equations after expanding the first one.

OpenStudy (anonymous):

@Jhannybean I am having a difficult time

OpenStudy (jhannybean):

Okay let's start you off. multiplying 2 to each term within the parenthesis, you get: -2x - 6y = 2 2x + 2y = 6 ________________ Did you see what i did?

OpenStudy (anonymous):

Ok yes! I know there is more to understand

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

I want to know how I solve the other one

OpenStudy (jhannybean):

The other one?

OpenStudy (anonymous):

Is still don't have a full understanding! @Jhannybean I have to solve more problems like this And I want to be able to do it on my own

OpenStudy (jhannybean):

Have you finished solving this one?

OpenStudy (anonymous):

No, I thought you were explaining the steps on how to solve it. So you got -2x-2y= 6 by multiplying 2 to each term. Idk what to do next or even where you got the number 2 :(

OpenStudy (jhannybean):

OK after the step i wrote above, you want to ADD the two equations together. -2x - 6y = 2 2x + 2y = 6 ___________________ 0x - 4y = 8 - 4y = 8 Divide both sides by -4 to isolate y, and you'll have your answer for the "y" portion

OpenStudy (anonymous):

so -4 divided by -4?That equals 1

OpenStudy (anonymous):

I am so confused @Jhannybean

OpenStudy (jhannybean):

yes thats correct,t hen what is 8/ -4 ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

I don't know the answer or the steps to get the answer. :( @Jhannybean

OpenStudy (jhannybean):

Okay, you're almost there!

OpenStudy (jhannybean):

Now that we have y, we can use this value to find x. Let's take the first equation for example. x + 3y = -1 If we use our value, y = -2, we can find x. x + 3(-2) = -1 x = -1 +6 x = ?

OpenStudy (anonymous):

5!!!!!!!!!!!!!!!!!! Holy cow the answer is D (5,-2) Could you please help me with a similar question. I don't get how you got the negative two :) @Jhannybean

OpenStudy (jhannybean):

We got the -2 when we eliminated x to solve for y. Look at the above posts.